OpenStudy (anonymous):

Simplify z/1 divide 1/7 A.1 B.z C.7 D.7z

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

oh sorry its z/7

OpenStudy (anonymous):

so would it be D?

OpenStudy (michele_laino):

we have to multiply the first fraction by the inverse of the second fraction, namely: \[\Large \frac{z}{1} \times \frac{7}{1} = ...?\]

OpenStudy (michele_laino):

so, it is option D

OpenStudy (anonymous):

yea but the z/1 is actually z/7

OpenStudy (michele_laino):

then we can write this: \[\Large \frac{z}{7} \times \frac{7}{1} = ...?\]

OpenStudy (anonymous):

yes but how you do that whith the numbers??

OpenStudy (anonymous):

how do you make it a fraction like that?

OpenStudy (michele_laino):

hint: